# On the spectral norm of Rademacher matrices

### v1

@@ 4. Proof of Theorem 5
If $l_{N}$ is the number of neighbours of $i_{N}$ in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N-1}\}^{\prime}$, then $l_{1}\geq l_{2}\geq\ldots\geq l_{M}$, so $Ml_{M}\leq|J_{k+r}(t)|$. Hence, using this algorithm we may find a subset $I\subset I_{k}(s)$ with cardinality $|I|\leq d_{A}^{-(r+18)/40}|J_{k+r}(t)|\leq d_{A}^{(r-13)/40}m$ such that for every $i\in I_{k}(s)\setminus I$, |$|\{j\in J_{k+r}(t)\setminus I^{\prime}\colon i\sim_{A}j\}|\leq\lceil d_{A}^{(r+18)/40}\rceil\leq 2d_{A}^{(r+18)/40}$. Note that if $(i,j)\in E_{A}$ and $(i,j)\in(I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, then $j\in J_{k+r}(t)\setminus I^{\prime}$. Therefore,
$\sum_{(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})}|a_{i,j}||s_{i}t_{j}|$ $\leq 2\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{(r+18)/40}d_{A}^{(1-k-r)/40}$
$\leq 2\|(a_{i,j})\|_{\infty}d_{A}^{19/40}\sum_{i\in I_{k}(s)}s_{i}^{2}.$ (14)
Observe that if $d_{A}^{-(r+18)/40}|J_{k+r}(t)|>m$ we may take $I=I_{k}(s)$ and then $a_{i,j}=0$ for $(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, so estimate (14) is also valid in this case.
@@
$\sum_{(i,j)\in(I_{k}(s)\times J_{k+r}(t))\cap(I^{\prime\prime}\times I^{\prime})}\varepsilon_{i,j}a_{i,j}s_{i}t_{j}\leq Y_{m,r}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2},$ (15)
@@
Define $Y_{r}:=\max_{1\leq m\leq n}Y_{m,r}$. Estimates (14) and (15) yield
@@
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\Biggl{(}\sum_{k\geq 1}2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}s_{i}^{2}+{\mathbb{E}}Y_{r}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}\Biggr{)}$
$\leq 2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}\sup_{\|s\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}^{2}\Bigr{)}^{1/2}\sup_{\|t\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq 2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}.$
Therefore, to establish Theorem 5 it is enough to prove the following lemma.
Lemma 19.
@@
First we show a connected counterpart to Lemma 19.
Lemma 20.
@@
${\mathbb{E}}\max_{1\leq k\leq n}\max_{I\in\mathcal{I}_{4}(k)}\ \sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}$ $\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Let us first fix $k$ and $I\in\mathcal{I}_{4}(k)$. Then $|I^{\prime}|\leq d_{A}k$ and $|I^{\prime\prime}|\leq d_{A}^{2}k$. Proposition 12, applied with $(a_{i,j})=(a_{i,j})_{i\in I^{\prime\prime},j\in I^{\prime}}$, $n=|I^{\prime\prime}|$, $m=|I^{\prime}|$, and $b=d_{A}^{3/8}k^{-1/2}$ yields
@@
By Lemma 11, $|\mathcal{I}_{4}(k)|\leq n(4d_{A}^{4})^{k}\leq\max\{n^{2},d_{A}^{12k}\}$ (recall that we assume that $d_{A}\geq 3$). We have
@@
Hence, by Proposition 10,
${\mathbb{E}}\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}$ $\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Applying again Proposition 10 and observing that
@@
Let us fix $r\geq 0$, $I\subset V$ such that $|I|\leq d_{A}^{(r-13)/40}m$, $s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}$ and $t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}$. Let $I_{1},\ldots,I_{l}$ be 4-connected components of $I$. Then $\{I_{1}^{\prime},\ldots,I_{l}^{\prime}\}$ is a partition of $I^{\prime}$, $\{I_{1}^{\prime\prime},\ldots,I_{l}^{\prime\prime}\}$ is a partition of $I^{\prime\prime}$ and
$\sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}=\sum_{u=1}^{l}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$ (16)
@@
$U_{1}$ :=$:=\bigl{\{}1\leq u\leq l\colon\ \|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}},$
$U_{2}$ :=$:=\bigl{\{}1\leq u\leq l\colon\ \|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}}.$
@@
$\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$ (19)
Bounds (16)-(19) yield
@@
and the assertion follows by Lemma 20 and Proposition 18. ∎
Acknowledgements. Part of this work was carried out while the author was visiting the Hausdorff Research Institute for Mathematics, Univerity of Bonn. The hospitality of HIM and of the organizers of the program Synergies between modern probability, geometric analysis and stochastic geometry is gratefully acknowledged. The author would like to thank Marta Strzelecka for careful readings of various versions of the manuscript and numerous valuable comments and Ramon van Handel for pointing out the reference [6].

### v2

@@ 4. Bounds up to polylog factors
If $l_{N}$ is the number of neighbours of $i_{N}$ in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N-1}\}^{\prime}$, then $l_{1}\geq l_{2}\geq\ldots\geq l_{M}$, so $Ml_{M}\leq|J_{k+r}(t)|$. Choose $M\leq m$ to be the maximal integer so that $l_{M}\geq d_{A}^{(r+18)/40}$ and set $I:=\{i_{1},\ldots,i_{M}\}$. This way we construct a subset $I\subset I_{k}(s)$ with cardinality $|I|\leq d_{A}^{-(r+18)/40}|J_{k+r}(t)|\leq d_{A}^{(r-13)/40}m$ such that for every $i\in I_{k}(s)\setminus I$, $|\{j\in J_{k+r}(t)\setminus I^{\prime}\colon i\sim_{A}j\}|<d_{A}^{(r+18)/40}$. Note that if $(i,j)\in E_{A}$ and $(i,j)\in(I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, then $j\in J_{k+r}(t)\setminus I^{\prime}$. Therefore,
$\sum_{(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})}|a_{i,j}||s_{i}t_{j}|$ $\leq\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{(r+18)/40}d_{A}^{(1-k-r)/40}$
(5.2) $\leq\|(a_{i,j})\|_{\infty}d_{A}^{19/40}\sum_{i\in I_{k}(s)}s_{i}^{2}.$
@@
(5.3) $\sum_{(i,j)\in(I_{k}(s)\times J_{k+r}(t))\cap(I^{\prime\prime}\times I^{\prime})}\varepsilon_{i,j}a_{i,j}s_{i}t_{j}\leq Y_{m,r}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2},$
@@
Define $Y_{r}:=\max_{1\leq m\leq n}Y_{m,r}$. The main advantage of introducing the variables $Y_{r}$ is that in their definition, the suprema are taken over delocalized vectors from the sphere of dimension $|I|$ with all coordinates bounded essentially by $d_{A}^{C}|I|^{-1/2}$, and Proposition 3.3 can be used to estimate these quantities.
Estimates (5.2) and (5.3) yield
@@
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\Biggl{(}\sum_{k\geq 1}d_{A}^{19/40}\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}s_{i}^{2}+{\mathbb{E}}Y_{r}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}\Biggr{)}$
$\leq d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}\sup_{\|s\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}^{2}\Bigr{)}^{1/2}\sup_{\|t\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}.$
Therefore, to establish Theorem 1.6 it is enough to prove the following lemma.
Lemma 5.1.
@@
First we show a connected counterpart to Lemma 5.1.
Lemma 5.2.
@@
${\mathbb{E}}\max_{1\leq k\leq n}\max_{I\in\mathcal{I}_{4}(k)}\ \sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\$$\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Let us first fix $k$ and $I\in\mathcal{I}_{4}(k)$. Then $|I^{\prime}|\leq d_{A}k$ and $|I^{\prime\prime}|\leq d_{A}^{2}k$. Proposition 3.3, applied with $(a_{i,j})=(a_{i,j})_{i\in I^{\prime\prime},j\in I^{\prime}}$, $n=|I^{\prime\prime}|$, $m=|I^{\prime}|$, and $b=d_{A}^{3/8}k^{-1/2}$ yields
@@
By Lemma 3.2, $|\mathcal{I}_{4}(k)|\leq n(4d_{A}^{4})^{k}\leq\max\{n^{2},d_{A}^{12k}\}$ (recall that we assume that $d_{A}\geq 3$). We have
@@
Hence, by Proposition 3.1,
${\mathbb{E}}\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\$$\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
@@
Applying again Proposition 3.1 and observing that
@@
Let us fix $0\leq r\leq 40$, $I\subset V$ such that $|I|\leq d_{A}^{(r-13)/40}m$, $s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}$ and $t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}$. Let $I_{1},\ldots,I_{l}$ be 4-connected components of $I$. Then $\{I_{1}^{\prime},\ldots,I_{l}^{\prime}\}$ is a partition of $I^{\prime}$, $\{I_{1}^{\prime\prime},\ldots,I_{l}^{\prime\prime}\}$ is a partition of $I^{\prime\prime}$ and
(5.4) $\sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}=\sum_{u=1}^{l}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
@@
$U_{1}$ $:=\bigl{\{}1\leq u\leq l\colon\ \|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}},$
$U_{2}$ $:=\bigl{\{}1\leq u\leq l\colon\ \|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}}.$
@@
(5.7) $\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$
Bounds (5.4)-(5.7) yield
@@
and the assertion follows by Lemma 5.2 and Proposition 4.4. ∎
Acknowledgements. Part of this work was carried out while the author was visiting the Hausdorff Research Institute for Mathematics, Univerity of Bonn. The hospitality of HIM and of the organizers of the program Synergies between modern probability, geometric analysis and stochastic geometry is gratefully acknowledged. The author would like to thank Marta Strzelecka for careful readings of various versions of the manuscript and numerous valuable comments, Ramon van Handel for pointing out the reference [7] and an anonymous referee for many thoughtful comments and for suggesting most of the examples presented in the paper.